OpenStudy (anonymous):

Need some help on a couple of math problems and could use some help and explaining please :) 1. (3x^2)^2 2. (7y^4)^2 3. xy^3z^-1 over x^2yz 4. y^2z over x^2yz^-3

OpenStudy (anonymous):

do you know know overall how to do these problems?

OpenStudy (anonymous):

not really thats why i asked for someone to help explain them to me :)

OpenStudy (anonymous):

true true the first answer is (3x^2)^2

OpenStudy (anonymous):

now try to figure out how i got that, tell me what u thinkin

OpenStudy (anonymous):

3: \[\large \frac{xy^3z^{-1}}{x^2yz}\] maybe easiest if you bring the \(z^{-1}\) in to the denominator and start with\[\large \frac{xy^3}{x^2yz^2}\]

OpenStudy (anonymous):

How can the first answer be the question asked? Does it not need to be simplified

OpenStudy (anonymous):

sorry autocorrect, its 9x^4

OpenStudy (anonymous):

@alex1248191 that was the question \[(3x^2)^2\] it is not really a question, it is an expression but i think they want you to square each term and get \[3^2x^4=9x^4\]

OpenStudy (anonymous):

I get it. @daydreamer just multiply by the power of two for each one.

OpenStudy (anonymous):

exactly @satellite73

OpenStudy (anonymous):

so what you did is squared all of it? so 3 * 3 =9 and then 2 * 2=4 okay i get it! It just clicked!!

OpenStudy (anonymous):

Yay! (:

OpenStudy (anonymous):

you square the first number, the you add both of the exponents together

OpenStudy (anonymous):

you don't square the exponents, (that mistake is comkonly made) you add them together

OpenStudy (anonymous):

oh okay that is noted most defiantly :)

OpenStudy (anonymous):

Nice! thats very important

OpenStudy (anonymous):

\[3\sqrt{2}+\sqrt{8}\] could you help with this last one :) @alex1248191